# Directional Diffusion-Style Code Editing Pre-training

### v1

@@ V-A2 Code-only editing scenario
Table IV shows the statistical information of the baselines and DivoT5 on the code-only editing scenario. In this scenario, we select code refinement without natural language guidance as the specific task. Due to the significant overlap between the old code and the new code in this task, the BLEU-related metrics tend to be high. Therefore, we primarily focus on the EM metric to assess the model’s performance.
From Table IV, we can observe that the DivoT5-small model with 60M parameters performs comparably to all the baselines with 220M parameters on the small version of the code refinement benchmark. The DivoT5-base (220M) model outperforms the CodeT5-base (220M) and CodeReviewer (220M) by more than 6% in the EM metric. On the medium version of the code refinement benchmark, the DivoT5-base model still outperforms the CodeT5-base (220M) model by 2.5 percentage points in EM metric. However, due to the high BLEU-related metrics (around 88%) for all models on this dataset, the DivoT5-base (220M) model does not surpass all other models in BLEU-related scores. Due to the lack of sufficient domain knowledge and guidance from natural language on the direction of changes, billion-scale models typically generate similar code but struggle to produce an exact string match. The knowledge of edit directions and patterns required by these datasets often needs to be acquired through fine-tuning the training set.
@@
Code Refinement
Models Small Medium
Metrics EM BLEU C-BLEU EM BLEU C-BLEU
CodeT5-small (60M) 19.06 76.23 76.44 10.92 89.20 87.25
CodeT5-base (220M) 21.61 77.43 77.24 13.96 87.64 87.05
CodeT5-large (770M) 21.7 77.38 77.14 14.76 89.22 87.35
CodeT5+ (220M) 22.18 78.27 77.48 15.13 88.64 86.28
CoditT5 (220M) 18.75 77.59 76.99 13.00 89.45 86.99
CCT5 (220M) 17.87 77.02 76.39 9.75 89.06 86.59
CodeReviewer (220M) 21.66 77.50 77.44 12.61 88.92 86.52
DeepSeekCoder-FS (6.7B-instruct) 1.73 68.40 74.26 0.66 77.71 76.94
Meta-Llama-3.1-FS (8B-instruct) 1.32 39.11 58.44 0.58 59.93 74.56
ChatGPT-FS (gpt-3.5-turbo) 1.90 73.40 76.64 1.10 86.19 82.71
DivoT5-small (60M) 22.31 77.29 77.74 14.62 88.53 86.08
DivoT5-base (220M) 23.02 77.65 78.03 16.55 88.58 86.31
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ1: We conduct experiments on four datasets for two code editing scenarios. The experimental results demonstrate that the DivoT5-base model significantly outperforms existing models. For example, on the CodeReview dataset, the DivoT5-base model achieves an EM metric improvement of around 10 percentage points compared to the CodeT5-base model. This highlights the superior performance of DivoT5 in capturing code editing characteristics.
@@ V-B RQ2: Generalizability to Non-Editing Scenario
The code translation task is different from the code editing scenarios, and we use it to evaluate the generalization capability of DivoT5. The code translation involves mapping one programming language to another, requiring the model to have a deep understanding of at least one language to produce accurate and meaningful translations. In this task, we evaluate the performance of translating Java code to CSharp code and vice versa. Table V demonstrates that even in the code translation task outside the context of editing existing code, DivoT5-base still exhibits improvements over the existing baseline models. For example, in the task of translating CSharp code to Java code, DivoT5-base outperforms the CCT5 model by 6.9 percentage points in EM metric, surpasses the CodeReviewer model by 3.2 percentage points, and exceeds the CodeT5-base model by 4.5 percentage points. Similarly, in the task of translating Java code to CSharp code, DivoT5-base outperforms CodeT5+ (220M) by over 5% in the EM metric. By analyzing the generated results, we find that the translation in this dataset is not simply a direct translation of the original code but also requires corresponding modifications, such as adding the ‘override‘ keyword at specific positions, importing and calling packages for certain functions, or even splitting a single line of code. Besides, regarding our translation task, the high similarity between Java and CSharp makes the transformation resemble code evolution, where the model decides between copying tokens or generating new ones without drastically altering the structure. These modification patterns are embedded in the training set and are difficult to capture through a few-shot approach, leading to poor performance from billion-scale models. However, when considering translation tasks in specific scenarios, DivoT5 demonstrates an advantage in capturing these modification and translation patterns, performing better than other models.
@@
DivoT5-small (60M) 68.30 92.03 91.90 72.50 91.36 91.38
DivoT5-base (220M) 69.60 92.34 92.13 73.20 91.38 91.56
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ2: To demonstrate the generalization capability of DivoT5, we conduct experiments on the code translation task. The results show that DivoT5 significantly outperforms other models in this task. For example, in translating Java code to CSharp code, DivoT5-base improves the EM score by over 3 percentage points compared to the CodeT5+ model with the same scale.
@@
Refine-Small
@@
DivoT5-small (60M) 40.16 22.31 72.50
w/o $KSM_{ED}$ 35.45 21.75 71.20
w/o $RM_{ED}$ 37.02 21.54 69.80
w/o $EDAE_{ED}$ 35.22 20.99 68.60
w/o $EDR$ 37.31 21.68 70.20
w/o $ALL$ 34.75 20.84 68.20
@@ V-C RQ3: Effectiveness of Pre-training Tasks
To evaluate the effectiveness of the proposed pre-training tasks, we conduct an extensive ablation experiment to explore the contribution of each pre-training task. Specifically, we remove the pre-training task separately to explore the model’s performance. Then, we remove all four pre-training tasks to assess the overall impact of pre-training datasets on the model’s performance. In this setting, our pre-training is to predict the new code on top of the old code. Due to computational resource limitations, we conduct this ablation study on only the small version of DivoT5. Due to resource constraints, ablation experiments are conducted only on DivoT5’s small version. For space limit, we selected three datasets for the ablation studies. All results can be found in our package [41].
Table VI presents the model performance after removing each pre-training task. It can be observed that the performance of the model declines when any task is removed, indicating that each pre-training task has a positive influence on the overall effects of the model. For example, in the CodeReview dataset, removing the $KSM_{ED}$ task (i.e., w/o $KSM_{ED}$) results in a decrease in the EM metric from 40.16% to 35.45%, while removing the $EDR$ task (i.e., w/o $EDR$) leads to a drop to 37.31%. In the small version of the code refinement dataset and the CSharp to Java code translation dataset, removing each task also causes significant performance degradation. For example, removing the $DAE_{ED}$ task reduces the EM metric on the code refinement dataset from 22.31% to 20.99%, and removing the RM task reduces the EM metric on the translation task from 72.50% to 69.80%. It is worth noting that, compared to removing other pre-training tasks, the removal of $KSM_{ED}$ task has a significant impact on the automated code review task and has the least impact on the code translation task. This is reasonable, as the $KSM_{ED}$ task focuses on fine-grained code editing, while the code translation task is distant from the editing scenario, thus suffering from the least impact. These results indicate that each pre-training task impacts the final performance.
@@
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Answer to RQ3: We compare the effect of removing a specific pre-training task from the DivoT5-small model on the EM metric. We find that the performance of the model decreases to varying degrees on three representative downstream datasets. This indicates that each pre-training task contributes to the final experimental performance.
@@ VI Discussion
Cost of the Directional Diffusion-Style Model. DivoT5, inspired by the gradual denoising concept in diffusion, treats old code during the evolution process as a form of noise and considers the new code as the final noise-free state to simulate real-world code editing scenarios. Instead of traditional diffusion, DivoT5 adopts an autoregressive approach, making its training and inference costs the same as encoder-decoder text generation models like Transformer and CodeT5 [27, 9]. To analyze the training and inference costs, let the sequence length be $L$ (i.e., $L_{\text{enc}}$ indicates the input length of the encoder and $L_{\text{dec}}$ indicates the output length of the decoder), the model’s hidden layer size be $H$, and the number of denoising steps in traditional diffusion is $T$. Typically, $L_{\text{enc}}$ and $L_{\text{dec}}$ are on the same order of magnitude, so both can be approximated as $L$. Thus, the training and inference costs of traditional diffusion are $O(TL^{2}H)$, where $L^{2}$ reflects interactions between $L_{\text{enc}}$ and $L_{\text{dec}}$ in the cross-attention mechanism. For an encoder-decoder text generation model with the same structure, the training phase has a time complexity of $O(L_{\text{enc}}^{2}H)$ for the encoder, $O(L_{\text{dec}}^{2}H)$ for the decoder, and $O(L_{\text{enc}}L_{\text{dec}}H)$ for the cross-attention layer. These result in an complexity of $O(L_{\text{enc}}^{2}H+L_{\text{enc}}L_{\text{dec}}H+L_{\text{dec}}^{2}H)$, yielding an overall time complexity of $O(L^{2}H)$. In the inference phase, due to the gradual accumulation of the output length, the time complexity increases to $O(L^{3}H)$.
Above all, the time complexity of the encoder-decoder model during the training phase ($O(L^{2}H)$) is lower than that of traditional diffusion ($O(TL^{2}H)$). In the inference phase, $T$ is typically much larger than the sequence length $L$. For example, to generate text within 200 tokens ($L$), $T$ may need to be set to 2000 [14, 15]. Consequently, $O(L^{3}H)$ is often significantly smaller than $O(TL^{2}H)$. In summary, the training and inference costs of DivoT5 are generally lower than those of traditional diffusion.
@@ VII Threats to Validity
Threats to internal validity might come from the implementation of our experiments. To mitigate this threat, we utilize open-source models available on HuggingFace [45] to minimize potential errors. Additionally, we fix the random seed to ensure experiment reproducibility.
Threats to external validity may arise from two factors. The first factor is the selection of the base model. To mitigate this potential threat, we choose a popular and representative model, CodeT5, as the base model. CodeT5 is widely recognized as a classic encoder-decoder framework and has been extensively applied in various tasks [9]. The second factor lies in downstream datasets used in our experiment. To mitigate this threat, we choose the widely used datasets from CodeXGLEU [16] and other existing code evolution models [8, 7]. To further generalize our results, we need to evaluate DivoT5 on more datasets. Since DivoT5 focuses on the gradual evolution from code to code, considering both copying and generating tokens, its performance improvement is minimal on tasks involving the transformation between massively different input and output, such as natural language to code or vice versa.
Threat to construct validity may arise from two factors. The first is the selection of evaluation metrics. To mitigate this concern, we choose widely used evaluation metrics in the field of code generation to assess the evolved code. For the automated code review task, we also select the same evaluation metrics as other code evolution models [8]. The second is related to data leakage. To mitigate this threat, we exclude data that overlaps with downstream tasks during the pre-training process. Specifically, we perform precise deduplication using a string-matching approach. If a sample from the downstream task’s test set appears in any pretraining data sample ‘x‘, we filter ‘x‘ out. Additionally, we conduct extensive ablation experiments to validate the effectiveness of our technical contributions.

### v2

@@ V-A2 Code-only editing scenario
Table IV summarizes the performance of DivoT5 and various baselines in the code-only editing scenario, which is evaluated across three representative datasets: Refine-S, Refine-M, and CodeReview-CodeOnly. In this setting, models are required to refine code without access to natural language guidance. Due to the substantial lexical overlap between old and new code versions, BLEU scores tend to be uniformly high and less discriminative. Therefore, we primarily emphasize the EM and SARI metrics, which more directly reflect the model’s ability to perform accurate and meaningful edits.
On the Refine-S dataset, DivoT5-small (60M) performs on par with or better than all 220M-parameter baselines, achieving 22.31 in EM and 47.38 in SARI. The DivoT5-base model (220M) achieves the highest overall performance, with 23.02 in EM and 47.92 in SARI, surpassing CodeT5-base and CodeReviewer by over 6% in exact match. These results demonstrate that DivoT5 learns effective editing strategies even without natural language cues, outperforming larger models with fewer parameters.
On the more challenging Refine-M dataset, which involves longer and more complex code segments, DivoT5-base continues to lead with 16.55 in EM and 43.09 in SARI, improving upon CodeT5-base by 2.59 percentage points in EM. While BLEU scores remain high (around 88) for all models due to large unchanged regions, DivoT5 still achieves competitive fluency (88.58 in BLEU), showing no trade-off between correctness and surface similarity.
On the CodeReview-CodeOnly dataset, which is constructed by removing the natural language guidance from the original CodeReview dataset, DivoT5 demonstrates even more pronounced advantages. Specifically, DivoT5-base achieves a 34.3% relative improvement in EM and a 10.9% relative improvement in SARI compared to CodeReviewer. In comparison to CodeT5+, DivoT5-base also shows clear gains, with EM increasing from 12.75 to 20.90 and SARI improving from 39.42 to 46.85. These results highlight the effectiveness of DivoT5 in code-only editing scenarios, particularly in the absence of external natural language signals.
Due to the lack of sufficient domain knowledge and guidance from natural language on the direction of changes, billion-scale few-shot models such as ChatGPT, Meta-Llama-3.1, and DeepSeekCoder perform poorly across all datasets. Although they often achieve reasonable BLEU scores by generating syntactically similar code, they fail to produce exact outputs or apply semantically faithful edits. The knowledge of edit directions and patterns required by these datasets often needs to be acquired through fine-tuning the training set.
@@
Models Refine-S Refine-M CodeReview-CodeOnly
Metrics EM BLEU SARI EM BLEU SARI EM BLEU SARI
CodeT5-small (60M) 19.06 76.23 44.69 10.92 89.20 38.73 10.77 82.18α 38.27
CodeT5-base (220M) 21.61 77.43α 45.78 13.96 87.64 40.25 13.15 82.56α 39.54
CodeT5-large (770M) 21.7 77.38α 46.03 14.76 89.22 41.08 13.34 82.58α 39.87
CodeT5+ (220M) 22.18 78.27 46.87 15.13 88.64α 42.40 12.75 82.67α 39.42
CoditT5 (220M) 18.75 77.59α 44.61 13.00 89.45 40.15 14.09 83.13 41.04
CCT5 (220M) 17.87 77.02 43.25 9.75 89.06 38.21 13.56 82.31α 41.59
CodeReviewer (220M) 21.66 77.50α 45.89 12.61 88.92α 39.57 15.56 81.97 42.26
CodeLlama (6.7B-base) 20.17 77.24α 45.78 13.23 88.30α 40.96 14.90 80.52 43.59
DeepSeekCoder-FS (6.7B-instruct) 1.73 68.40 35.25 0.66 77.71 33.41 11.03 73.59 40.24
Meta-Llama-3.1-FS (8B-instruct) 1.32 39.11 35.34 0.58 59.93 33.33 12.14 74.28 40.37
ChatGPT-FS (gpt-3.5-turbo) 1.90 73.40 37.29 1.10 86.19 36.56 13.52 77.30 42.51
DivoT5-small (60M) 22.31 77.29α 47.38 14.62 88.53α 42.37 17.00 82.68α 43.71
DivoT5-base (220M) 23.02 77.65α 47.92 16.55 88.58α 43.09 20.90 82.47α 46.85
Answer to RQ1: We conduct experiments on seven datasets for two code editing scenarios. The experimental results demonstrate that the DivoT5-base model significantly outperforms existing models. For example, on the CodeReview dataset, the DivoT5-base model achieves an EM metric improvement of around 10 percentage points compared to the CodeT5-base model. This highlights the superior performance of DivoT5 in capturing code editing characteristics.
@@ V-B RQ2: Generalizability to Non-Editing Scenario
To assess the generalization capability of DivoT5 beyond code editing, we introduce the code translation task—a distinct scenario that requires the model to translate code from one programming language to another. In this task, we evaluate the performance of translating Java code to CSharp code and vice versa. Table V demonstrates that even in the code translation task outside the context of editing existing code, DivoT5-base still exhibits improvements over the existing baseline models. For example, in the task of translating CSharp code to Java code, DivoT5-base outperforms the CCT5 model by 6.9 percentage points in EM metric, surpasses the CodeReviewer model by 3.2 percentage points, and exceeds the CodeT5-base model by 4.5 percentage points. Similarly, in the task of translating Java code to CSharp code, DivoT5-base outperforms CodeT5+ (220M) by over 5% in the EM metric. By analyzing the generated results, we find that the translation in this dataset is not simply a direct translation of the original code but also requires corresponding modifications, such as adding the ‘override‘ keyword at specific positions, importing and calling packages for certain functions, or even splitting a single line of code. Besides, regarding our translation task, the high similarity between Java and CSharp makes the transformation resemble code evolution, where the model decides between copying tokens or generating new ones without drastically altering the structure. These modification patterns are embedded in the training set and are difficult to capture through a few-shot approach, leading to poor performance from billion-scale models. However, when considering translation tasks in specific scenarios, DivoT5 demonstrates an advantage in capturing these modification and translation patterns, performing better than other models.
@@
DivoT5-small (60M) 68.30 92.03 91.90 72.50 91.36α 91.38
DivoT5-base (220M) 69.60 92.34α 92.13 73.20 91.38α 91.56
Answer to RQ2: To demonstrate the generalization capability of DivoT5, we conduct experiments on the code translation task. The results show that DivoT5 significantly outperforms other models in this task. For example, in translating Java code to CSharp code, DivoT5-base improves the EM score by over 3 percentage points compared to the CodeT5+ model with the same scale.
@@
NL-CodeRefinement
EM
BugFixing-S
EM
BugFixing-M
EM
Refine-S
EM
Refine-S
EM
CodeReview-CodeOnly
EM
Java-CSharp
@@
[…]
w/o $KSM_{ED}$ 35.45 28.69 35.28 27.36 21.75 13.67 14.44 68.00 71.20
w/o $RM_{ED}$ 37.02 28.61 35.30 28.10 21.54 13.91 13.68 66.30 69.80
w/o $EDAE_{ED}$ 35.22 28.74 35.50 27.65 20.99 13.86 11.76 67.50 68.60
w/o EDREDR 37.31 28.72 34.76 28.37 21.68 14.03 15.83 67.50 70.20
w/o ALLALL 34.75 28.50 35.35 27.36 20.84 13.41 12.05 67.60 68.20
@@ V-C RQ3: Effectiveness of Pre-training Tasks
To evaluate the effectiveness of the proposed pre-training tasks, we conduct an extensive ablation experiment to explore the contribution of each pre-training task. Specifically, we remove the pre-training task separately to explore the model’s performance. Then, we remove all four pre-training tasks to assess the overall impact of pre-training datasets on the model’s performance. In this setting, our pre-training is to predict the new code on top of the old code. Due to computational resource limitations, we conduct this ablation study on only the small version of DivoT5. Due to resource constraints, these ablation studies are conducted on the 60M-parameter variant of DivoT5, which provides an efficient setting for experimentation under limited computational resources. This choice is consistent with prior work such as DeepSeekCoder [guo2024deepseekcoder], where Fill-in-the-Middle (FIM) ablation experiments were conducted on the 1.3B model—nearly 30× smaller than their largest 33B variant—demonstrating a practical strategy for resource-efficient experimentation. Furthermore, our approach follows established practices in the CodeT5 model family [codet52021, zhu2024grammart5], where 60M-parameter variants are frequently used to examine the effects of different pre-training strategies. Since DivoT5 is also built on the CodeT5 architecture, employing a 60M model for ablation ensures both computational tractability and methodological consistency with previous studies.
[…]
@@
Answer to RQ3: We compare the effect of removing a specific pre-training task from the DivoT5-small model on the EM metric. We find that the performance of the model decreases to varying degrees on three representative downstream datasets. This indicates that each pre-training task contributes to the final experimental performance.
@@ VI Discussion
Cost of the Directional Diffusion-Style Model. DivoT5, inspired by the gradual denoising concept in diffusion, treats old code during the evolution process as a form of noise and considers the new code as the final noise-free state to simulate real-world code editing scenarios. Instead of traditional diffusion, DivoT5 adopts an autoregressive approach, making its training and inference costs the same as encoder-decoder text generation models like Transformer and CodeT5 [transformer2017, codet52021]. To analyze the training and inference costs, let the sequence length be $L$ (i.e., $L_{\text{enc}}$ indicates the input length of the encoder and $L_{\text{dec}}$ indicates the output length of the decoder), the model’s hidden layer size be $H$, and the number of denoising steps in traditional diffusion is $T$. Typically, $L_{\text{enc}}$ and $L_{\text{dec}}$ are on the same order of magnitude, so both can be approximated as $L$. Thus, the training and inference costs of traditional diffusion are $O(TL^{2}H)$, where $L^{2}$ reflects interactions between $L_{\text{enc}}$ and $L_{\text{dec}}$ in the cross-attention mechanism. For an encoder-decoder text generation model with the same structure, the training phase has a time complexity of $O(L_{\text{enc}}^{2}H)$ for the encoder, $O(L_{\text{dec}}^{2}H)$ for the decoder, and $O(L_{\text{enc}}L_{\text{dec}}H)$ for the cross-attention layer. These result in an complexity of $O(L_{\text{enc}}^{2}H+L_{\text{enc}}L_{\text{dec}}H+L_{\text{dec}}^{2}H)$, yielding an overall time complexity of $O(L^{2}H)$. In the inference phase, due to the gradual accumulation of the output length, the time complexity increases to $O(L^{3}H)$.
Above all, the time complexity of the encoder-decoder model during the training phase ($O(L^{2}H)$) is lower than that of traditional diffusion ($O(TL^{2}H)$). In the inference phase, $T$ is typically much larger than the sequence length $L$. For example, to generate text within 200 tokens ($L$), $T$ may need to be set to 2000 [difflm_ARDiff_wu2024ar, difflm_DiffLM_li2022diffusion]. Consequently, $O(L^{3}H)$ is often significantly smaller than $O(TL^{2}H)$. In summary, the training and inference costs of DivoT5 are generally lower than those of traditional diffusion, and can therefore also be effectively adopted for scaling to larger models or alternative architectures.
@@
Generalization and Extensibility. DivoT5 is designed with strong generalization potential across both code granularity levels and model architectures. While it is trained primarily on token-level code edits, this level of granularity serves as the foundation for modeling more complex transformations. In real-world software maintenance, many structural changes, such as line-level updates, block replacements, or module-level refactorings, are composed of sequences of fine-grained edits. By modeling token-level operations, DivoT5 captures the fundamental building blocks of such transformations and has potential extensions toward hierarchical or semantic-aware edit modeling. Recent studies [edit_emperical, edit_emperical_2] also emphasize the prevalence and practical significance of token-level modifications in real development workflows, further validating our design choice. In addition, DivoT5 aims to simulate real-world developer behavior by leveraging the intermediate states of code evolution as training signals. Instead of relying solely on before-and-after code pairs, we extract supervision from the actual editing trajectory, which more closely reflects how developers iteratively modify code. These evolution directions are architecture-agnostic and can be applied to various model types. Although we use an encoder-decoder architecture in this work, the same training data and objectives can be easily adapted to other architectures (e.g., decoder-only architecture) or larger language models. This design makes our approach flexible, extensible, and compatible with future model developments.
@@ VII Threats to Validity
Threats to internal validity might come from the implementation of our experiments. To mitigate this threat, we utilize open-source models available on HuggingFace [huggingface] to minimize potential errors. Additionally, we fix the random seed to ensure experiment reproducibility.
Threats to external validity may arise from two factors. The first factor is the selection of the base model. To mitigate this potential threat, we choose a popular and representative model, CodeT5, as the base model. CodeT5 is widely recognized as a classic encoder-decoder framework and has been extensively applied in various tasks [codet52021]. The second factor lies in the downstream datasets used in our experiment. To mitigate this threat, we choose the widely used datasets from CodeXGLEU [codexglue2021] and other existing code evolution models [coditt52022, codereviewer]. To further generalize our results, we need to evaluate DivoT5 on more datasets. Nonetheless, we acknowledge that important code editing tasks, such as code refactoring, which improves structure without changing functionality, are not covered in this work. In future work, we plan to scale DivoT5 to larger models and extend it to broader scenarios, including refactoring and architectural changes. As DivoT5 is designed for gradual code evolution with both copying and generation, its advantages are less evident in tasks with large input-output gaps, such as natural language to code or vice versa.
Threat to construct validity may arise from two factors. The first is the selection of evaluation metrics. To mitigate this concern, we choose widely used evaluation metrics in the field of code generation to assess the evolved code. For the CodeReview dataset, we also select the same evaluation metrics as other code evolution models [coditt52022]. The second is related to data leakage. To mitigate this threat, we exclude data that overlaps with downstream tasks during the pre-training process. Specifically, we perform precise deduplication using a string-matching approach. If a sample from the downstream task’s test set appears in any pretraining data sample ‘x’, we filter ‘x’ out. Additionally, we conduct extensive ablation experiments to validate the effectiveness of our technical contributions.